# Chance-Constrained Covariance Steering for Discrete-Time Markov Jump Linear Systems

### v3

@@ IV-B Iterative Convex Control Synthesis Framework
Using the deterministic formulation of chance constraints in Lemmas 2 and 3, we can now incorporate them in our two-step optimization framework. Note that the chance constraints introduce two-way coupling in the problem, while previously only the covariance problem depended on the mean trajectory and not vice-versa. To solve this, we propose a hierarchical optimization framework which iteratively solves the mean and covariance optimization problem by relaxing the chance constraints. We relax the constraints by adding a slack variable, and then penalize its 1-norm in the objective function. Using the 1-norm promotes sparsity of the slack variable, and has demonstrated favorable convergence properties in other hierarchical optimization frameworks [27, 28]. We now define the relaxed mean and covariance problems as follows:
@@
[…]
$a_{j}^{\top}\mu_{k}$ $+b_{j}+m_{\mathcal{N}}(\delta_{x,j})\sqrt{a_{j}^{\top}\Sigma_{k}a_{j}}\leq\beta_{1,k}$ (29c)
$p_{j}^{\top}\overline{u}_{k}(i)$ $+q_{j}+m_{\mathcal{N}}(\delta_{u,j}(i))\sqrt{p_{j}^{\top}Y_{k}(i)p_{j}}\leq\beta_{2,k}$ (29d)
$\|\overline{u}_{k}(i)\|+$ $m_{\chi^{2}}(\epsilon_{u}(i),n_{u})\sqrt{\lambda_{\max}(Y_{k}(i))}\leq u_{\max}(i)+\beta_{3,k}$ (29e)
@@
[…]
$m^{2}_{\mathcal{N}}(\delta_{x,j}){a_{j}^{\top}\Sigma_{k}a_{j}}-(a_{j}^{\top}\mu_{k}+b_{j})^{2}$ $-\zeta_{1,j}\leq 0,$ (30c)
$m^{2}_{\mathcal{N}}(\delta_{u,j}){p_{j}^{\top}Y_{k}(i)p_{j}}-(p_{j}^{\top}\mu_{k}+d_{j})^{2}$ $-\zeta_{2,j}\leq 0,$ (30d)
$m^{2}_{\chi^{2}}(\epsilon_{x},n_{x})\lambda_{\max}(\Sigma_{k})$ $-\zeta_{3,j}\leq 0$ (30e)
$m^{2}_{\chi^{2}}(\epsilon_{u}(i),n_{u})\lambda_{\max}(Y_{k}(i))-u_{m}^{2}$ $-\zeta_{4,j}\leq 0$ (30f)
@@
Under Assumptions 1, 2 and 4, the optimal solution to Problem 7 satisfies $Y_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$.
@@
$+\sum_{i\in\Omega}\left\{\mathrm{Tr}(M^{\top}_{k}(i)C_{k}(i))+\lambda_{u,n}(i)e_{u,n}(i)\right.$
@@
where $C_{k}(i)$ is given in eq. 15 and $e_{x,h}$, $e_{u,h}$, $e_{x,n}$, and $e_{u,n}$ represent eqs. 30c, 30d, 30e and 30f respectively with $\lambda_{x,h}$, $\lambda_{u,h}$, $\lambda_{x,n}$, and $\lambda_{u,n}$ being the respective Lagrange multipliers. $\Lambda_{k}$,$M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively, and are all symmetric as explained in the proof of theorem 1. The relevant first-order KKT conditions for optimality are:
$\tfrac{\partial\mathcal{L}}{\partial L_{k}(i)}=2\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}A_{k}(i)+2M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}=0$ (32a)
$\tfrac{\partial\mathcal{L}}{\partial Y_{k}(i)}=\rho_{k}(i)R_{k}-M_{k}(i)+\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u}(i),n_{u})v_{k}(i)v_{k}^{\top}(i)$
$+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}+\sum_{j\in\Omega}\pi^{(k)}_{ij}B^{\top}_{k}(i)\Lambda_{k}B_{k}(i)=0$ (32b)
@@
where $\pi^{(k)}_{ij}=p_{ij}\rho_{k}(i)$ and $v_{k}(i)$ is the eigenvector of $Y_{k}(i)$ corresponding to its maximum eigenvalue. We now prove by contradiction that $C_{k}(i)=0$, or the relaxation is lossless. Let us suppose that $C_{k}(i)$ has at least one nonzero eigenvalue, which implies $M_{k}(i)$ is singular from Lemma 1 and eq. 32c. Rearranging eq. 32a, we get $\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}A_{k}^{-1}(i)$. Using this and the substitutions defined in eq. 12, we simplify eq. 32b as: $\rho_{k}(i)R_{k}+\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u},n_{u})v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$. Representing the LHS as $\Delta_{k}(i)$ and taking the determinant of both sides, we get:
$\mathrm{det}(\Delta_{k}(i))=\mathrm{det}(M_{k}(i))\mathrm{det}(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$
From Assumption 1, we have that $\rho_{k}(i)R_{k}\succ 0$. Now we know $\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u},n_{u})v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}\geq 0$ which implies $\Delta_{k}(i)>0$, resulting in a contradiction. Hence, $C_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)-Y_{k}(i)=0\;\forall k\in\mathbb{Z}_{0}^{T-1},i\in\Omega$. ∎
@@ IV-C Sequential Convex Optimization Framework
7:     if $\max_{m}\|\beta_{m}\|_{\infty}$$>$tol or $\max_{m}\|\zeta_{m}\|_{\infty}$ $>$tol then
@@ V Numerical Simulations
$A_{k}(1)$ $=\begin{bmatrix}-0.2&1\\ -0.1&0.1\end{bmatrix}$ $B_{k}(1)$ $=\begin{bmatrix}1&0\\ 2&0\end{bmatrix}$ $G_{k}(1)$ $=\begin{bmatrix}0.1&0\\ 0&0.1\end{bmatrix}$
@@
$\mathbb{P}$ $=\begin{bmatrix}0.7&0.3\\ 0.9&0.1\end{bmatrix}$ $c_{k}$ $=\begin{bmatrix}0.01\\ 0.01\end{bmatrix}$ $T$ $=6$
The initial and final conditions are set as $\rho_{0}=[0.5,0.5]$, $\overline{\mu}_{0}=[40,10]^{\top}$, $\overline{\Sigma}_{0}=6I$ and $\mu_{f}=[5,10]^{\top}$, $\Sigma_{f}=2.5I$. We set a half-space state chance constraint of $P(a^{\top}x>b)\geq 0.95$, where $a=[1,0]^{\top}$ and $b=-10$ and a control norm bound constraint $P(\|u_{k}(i)\|\leq u_{\max})\geq 0.95$ where $u_{\max}=5$.
We use the CVXPY package[29] with MOSEK solver[30] to run Algorithm 1 on a 2023 Macbook Pro M1 Pro laptop with 16 GB of RAM. All initial weights are set at $\alpha_{m}=1e2$, with $\eta=5$ and $tol=1e-6$. It takes 3.43 seconds and $3$ iterations for Algorithm 1 to converge to the preset tolerance values. Figures 1(a) and 1(b) plot 1000 Monte Carlo trajectories for this problem without and with state chance constraints, respectively. Both figures use the control norm chance constraint. The initial covariance, terminal covariance, and the mean trajectory are shown in green, red, and blue, respectively, while Monte Carlo trajectories are shown in grey. The unshaded area indicates the feasible region under chance constraints, enforced using the state half-space constraint. Figure 2 shows the control norm history for the Monte Carlo simulations. None of the samples disobeyed the state chance constraints, while the control norm exceeded $u_{\max}$ in $1.3\%$ of the total sample size, which meets the prescribed risk bound of $5\%$. These numerical results demonstrate the performance of the controller to steer the covariance within the final target value in the given time horizon, while respecting the chance constraints. Moreover, the state trajectories in Figure 1 demonstrate the multi-modal nature of the problem because of the Markovian jump dynamics and mode-dependent controllers. We also note the number of state distribution modes increasing over time, most apparent in Figure 1(a), clearly demonstrating the exponential Gaussian mixture distribution of MJLS, which is hard to control using conventional stochastic optimal control approaches.
@@
Figure 1: State trajectories for Monte Carlo simulations with the state chance constraint $P(x_{1}\geq-10)\geq 0.95$ (shaded region: infeasible.) Both cases use $P(\|u_{k}(i)\|\leq 5)\leq 0.95$.
Figure 2: Control Norm History for the Monte Carlo simulations with state chance constraint $P(x_{1}\geq-10)\geq 0.95$.
@@ VI Conclusion
In this paper, we consider the optimal covariance steering problem for MJLS under chance constraints. We demonstrate the coupled nature of the problem and decompose the problem into a mean-covariance problem, which can be solved in a two-step convex optimization framework. We propose lossless convex relaxations of the covariance subproblem in both unconstrained and chance-constrained cases and discuss an iterative convex optimization framework to solve the chance-constrained covariance steering problem. Numerical simulations for the covariance steering framework are presented. Future work will analyze convergence behavior of the framework and extensions to semi-markov jump systems.
@@ Appendix A MJLS identities
$\mathbb{E}[x_{k}u_{k}^{\top}(i)\mathds{1}_{r_{k}=i}]=\mathbb{E}[x_{k}(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))^{\top}\mathds{1}_{r_{k}=i}]$
$=\rho_{k}(i)\left[\mu_{k}\overline{u}^{\top}_{k}(i)+{\Sigma}_{k}K^{\top}_{k}(i)\right]$ (33)
$\mathbb{E}[u_{k}(i)c_{k}^{\top}\mathds{1}_{r_{k}=i}]=\mathbb{E}[(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))c_{k}^{\top}\mathds{1}_{r_{k}=i}]$
@@
$\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]=\mathbb{E}[(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))(\overline{u}_{k}(i)+$
$\hskip 142.26378ptK_{k}(i)(x_{k}-\mu_{k}))^{\top}\mathds{1}_{r_{k}=i}]$
$=\mathbb{E}[\mathds{1}_{r_{k}=i}]\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+\overline{u}_{k}(i)(\mathbb{E}[x^{\top}_{k}\mathds{1}_{r_{k}=i}]-$
$\mu^{\top}_{k}\mathbb{E}[\mathds{1}_{r_{k}=i}])K_{k}^{\top}(i)+K_{k}(i)(\mathbb{E}[x_{k}\mathds{1}_{r_{k}=i}]-\mu_{k}\mathbb{E}[\mathds{1}_{r_{k}=i}])\overline{u}_{k}^{\top}(i)$
$\hskip 42.67912pt+K_{k}(i)\mathbb{E}[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\mathds{1}_{r_{k}=i}]K_{k}^{\top}(i)$
$=\rho_{k}(i)\left[\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+K_{k}(i){\Sigma}_{k}K_{k}^{\top}(i)\right]$ (35)
Appendix B Discussion on independence of $x_{k}$ and $r_{k}$
In this section, we provide numerical results showing that the relationship between random variables $x_{k}$ and $r_{k}$ is neglible and they can therefore, be considered independent for all practical purposes.
We first compare $q_{k}(i)$ and $\rho_{k}(i)\mu_{k}$. A close agreement between the two would indicate our initial assumption. We consider a scalar uncontrolled ($B_{k}(i)=0$) MJLS system without noise or bias ($G_{k}(i)=c_{k}=0$) to simplify the analysis. The system parameters are given below:
$A_{k}(1)=0.8$ $\quad A_{k}(2)=1.2$ $\mathbb{P}=\begin{bmatrix}0.7&0.3\\ 0.9&0.1\end{bmatrix}$ $\quad\rho_{k}(i)=[0.7;0.3]$
Figure 3 shows three plots which compare the sample mean $\mu_{k}$ multiplied by $\rho_{k}(i)$ to $q_{k}(i)$ which is calculated analytically using eq. 4. The plots show high agreement between $q_{k}(i)$ and $\rho_{k}(i)\mu_{k}$ indicating that the dependence between $\mathds{1}_{r_{k}=i}$ and $x_{k}$ is neglible.
Figure 3: Figure shows comparison between $q_{k}(i)$ and $\rho_{k}(i)\mu_{k}$, where $\mu_{k}$ is calculated using samples and $q_{k}(i)$ is calculated analytically for both modes. The right most plot shows the comparison of sampled and analytical mean $\mu_{k}$, demonstrating that the number of samples considered are sufficient for statistical convergence.
To further demonstrate this, and also provide numerical evidence to prove our mean and covariance propagation equations are correct, we consider the following time-varying system:
$A_{1}(k)$ $=\begin{bmatrix}-0.5&1\\ -1&\sin(0.2\pi k)\end{bmatrix}$ $A_{2}(k)$ $=\begin{bmatrix}\cos(0.2\pi k)&0.5\\ 0.5&0.5\end{bmatrix}$ (36)
$B_{1}(k)$ $=\begin{bmatrix}1\\ 0\end{bmatrix}$ $B_{2}(k)$ $=\begin{bmatrix}0.5\\ 1\end{bmatrix}$
$G_{1}(k)$ $=\begin{bmatrix}1\\ 1.2\end{bmatrix}$ $G_{2}(k)$ $=\begin{bmatrix}0.2\\ 0.3\sin(0.2\pi k)\end{bmatrix}$
$T$ $=10$ $\mathbb{P}$ $=\begin{bmatrix}0.6&0.4\\ 0.5&0.5\end{bmatrix}$
We use a randomized open loop control law sampled uniformly from $[-10,10]$. The noise is $w_{k}\sim\mathcal{N}(1,1000)$ and the initial state distribution is given by $x_{0}\sim\mathcal{N}(\begin{bmatrix}20,100\end{bmatrix}^{\top},10I_{2})$. Figure 4 plots the comparison between sampled mean and covariance values with theoretical calculations using the mean and covariance propagation equations provided in this work for different number of samples. Our simulations demonstrate the numerical accuracy of the propagation equations, while also validating our assumption about $x_{k}$ and $r_{k}$.
Figure 4: Comparison of sampled mean and covariance values with theoretical values calculated using propagation equations for 10,100 and 1e5 samples.

### v4

@@ IV-B Iterative Convex Control Synthesis Framework
Using the deterministic formulation of chance constraints in Lemmas 2 and 3, we can now incorporate them in our two-step optimization framework. Note that the chance constraints introduce two-way coupling in the problem, while previously only the covariance problem depended on the mean trajectory and not vice-versa. To solve this, we propose a hierarchical optimization framework which iteratively solves the mean and covariance optimization problem by relaxing the chance constraints. We relax the constraints by adding a slack variable, and then penalize its 1-norm in the objective function. Using the 1-norm promotes sparsity of the slack variable, and has demonstrated favorable convergence properties in other hierarchical optimization frameworks [27, 28]. We now define the relaxed mean and covariance problems in Problem 6 and Problem 7 respectively as follows:
@@
$a_{j}^{\top}\mu_{k}$ $+b_{j}+\sqrt{\frac{1-\delta_{x,j}}{\delta_{x,j}}a_{j}^{\top}\Sigma_{k}a_{j}}\leq\beta_{1,k},$ (27c)
$f_{j}^{\top}\overline{u}_{k}(i)$ $+g_{j}+\sqrt{\frac{1-\delta_{u,j}(i)}{\delta_{u,j}(i)}p_{j}^{\top}Y_{k}(i)p_{j}}\leq\beta_{2,k}$ (27d)
$\|\overline{u}_{k}(i)\|+$ $\sqrt{\frac{n_{u}(i)}{\epsilon_{u}(i)}\lambda_{\max}(Y_{k}(i))}\leq u_{\max}(i)+\beta_{3,k}$ (27e)
@@
${\frac{1-\delta_{x,j}}{\delta_{x,j}}a_{j}^{\top}\Sigma_{k}a_{j}}-(a_{j}^{\top}\mu_{k}+b_{j})^{2}$ $\leq\zeta_{1,j},$ (28c)
${\frac{1-\delta_{u,j}(i)}{\delta_{u,j}(i)}p_{j}^{\top}Y_{k}(i)p_{j}}-(f_{j}^{\top}\mu_{k}+g_{j})^{2}$ $\leq\zeta_{2,j},$ (28d)
$\frac{n_{x}}{\epsilon_{x}}\lambda_{\max}(\Sigma_{k})$ $\leq\zeta_{3,j}$ (28e)
$\frac{n_{u}(i)}{\epsilon_{u}(i)}\lambda_{\max}(Y_{k}(i))-u_{m}^{2}$ $\leq\zeta_{4,j}$ (28f)
@@
Under Assumptions 1, 2 and 4, the optimal solution to Problem 7 satisfies $Y_{k}(i)=L_{k}(i){{S}^{-1}_{k}(i)}L_{k}^{\top}(i)$.
@@
$+\sum_{i\in\mathcal{S}}\left\{\mathrm{Tr}(M^{\top}_{k}(i)C_{k}(i))+\lambda_{u,n}(i)e_{u,n}(i)\right$.
@@
where $C_{k}(i)$ is given in eq. 13 and $e_{x,h}$, $e_{u,h}$, $e_{x,n}$, and $e_{u,n}$ represent eqs. 28c, 28d, 28e and 28f respectively with $\lambda_{x,h}$, $\lambda_{u,h}$, $\lambda_{x,n}$, and $\lambda_{u,n}$ being the respective Lagrange multipliers. $\Lambda_{k}$$M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively, and are all symmetric as explained in the proof of theorem 1. The relevant first-order KKT conditions for optimality are:
$\tfrac{\partial\mathcal{L}}{\partial L_{k}(i)}=2\sum_{j\in\mathcal{S}}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}A_{k}(i)+2M_{k}(i)L_{k}(i)S_{k}^{-1}(i)$
$\hskip 199.16928pt=0$ (30a)
$\tfrac{\partial\mathcal{L}}{\partial Y_{k}(i)}=\rho_{k}(i)R_{k}-M_{k}(i)+\lambda_{u,n}\tfrac{n_{u}(i)}{\epsilon_{u}(i)}v_{k}(i)v_{k}^{\top}(i)$
$+\lambda_{u,h}\tfrac{1-\delta_{u,j}(i)}{\delta_{u,j}(i)}f_{j}f^{\top}_{j}+\sum_{j\in\mathcal{S}}\pi^{(k)}_{ij}B^{\top}_{k}(i)\Lambda_{k}B_{k}(i)=0$ (30b)
@@
where $\pi^{(k)}_{ij}=p_{ij}\rho_{k}(i)$ and $v_{k}(i)$ is the eigenvector of $Y_{k}(i)$ corresponding to its maximum eigenvalue. We now prove by contradiction that $C_{k}(i)=0$, or the relaxation is lossless. Let us suppose that $C_{k}(i)$ has at least one nonzero eigenvalue, which implies $M_{k}(i)$ is singular from Lemma 1 and eq. 30c. Rearranging eq. 30a, we get $\sum_{j\in\mathcal{S}}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i)S_{k}^{-1}(i)A_{k}^{-1}(i)$. Using this and substitutions defined in eq. 10, we simplify eq. 30b as: $\rho_{k}(i)R_{k}+\lambda_{u,n}\tfrac{n_{u}(i)}{\epsilon_{u}(i)}v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}\tfrac{1-\delta_{u,j}(i)}{\delta_{u,j}(i)}f_{j}f^{\top}_{j}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$. Representing the LHS as $\Delta_{k}(i)$ and taking the determinant of both sides, we get:
$\mathrm{det}(\Delta_{k}(i))=\mathrm{det}(M_{k}(i))\mathrm{det}(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))=0$
From Assumption 1, we have that $\rho_{k}(i)R_{k}\succ 0$. Now we know $\lambda_{u,n}\tfrac{n_{u}(i)}{\epsilon_{u}(i)}v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}\tfrac{1-\delta_{u,j}(i)}{\delta_{u,j}(i)}f_{j}f^{\top}_{j}\geq 0$ which implies $\Delta_{k}(i)>0$, resulting in a contradiction. Hence, $C_{k}(i)=L_{k}(i){S}^{-1}_{k}(i)L_{k}^{\top}(i)-Y_{k}(i)=0\;\forall k\in\mathbb{Z}_{0}^{T-1},i\in\mathcal{S}$. ∎
@@ IV-C Sequential Convex Optimization Framework
7:  if $\max_{m}\|\beta_{m}\|_{\infty}$>>tol or $\max_{m}\|\zeta_{m}\|_{\infty}$ >>tol then
@@ V Numerical Simulations
$A_{k}(1)$ $=\begin{bmatrix}-0.2&1\\ -0.1&0.1\end{bmatrix}$ $B_{k}(1)$ $=\begin{bmatrix}1&0.5\\ 2&0\end{bmatrix}$ $G_{k}(1)$ $=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$
@@
$\mathbb{P}$ $=\begin{bmatrix}0.8&0.2\\ 0.9&0.1\end{bmatrix}$ $c_{k}$ $=\begin{bmatrix}0.01\\ 0.01\end{bmatrix}$ $T$ $=6$
The initial and final conditions are set as $\rho_{0}=[0.3,0.7]$, ${\mu}_{0}=[25,40]^{\top}$, ${\Sigma}_{0}=6I$ and $\mu_{f}=[5,10]^{\top}$, $\Sigma_{f}=3I$. We set a half-space state chance constraint of $\mathbb{P}(a^{\top}x+b<0)\geq 0.95$, where $a=[0,-1]^{\top}$ and b=10b=-10 and a control norm bound constraint $\mathbb{P}(\|u_{k}(i)\|\leq u_{\max})\geq 0.95$ where $u_{\max}=8$. We consider a zero-mean gaussian distribution $\mathcal{N}(0,I_{n_{w}})$ for $w_{k}$. For sampling $x_{0}$, we also consider a normal distribution with mean $\mu_{0}$ and covariance $\Sigma_{0}$.
We use the CVXPY package[29] with MOSEK solver[30] to run Algorithm 1 on a 2023 Macbook Pro M1 Pro laptop with 16 GB of RAM. All initial weights are set at $\alpha_{m}=10^{2}$, with $\eta=1.5$ and $tol=10^{-6}$. It takes 3.43 seconds and $7$ iterations for Algorithm 1 to converge to the preset tolerance values. Figures 2(a) and 2(b) plot 2500 Monte Carlo trajectories for this problem without and with state chance constraints, respectively. In both cases, the control norm chance constraint is also imposed. The initial covariance, terminal covariance, and the mean trajectory are shown in green, red, and blue, respectively, while Monte Carlo trajectories are shown in grey. The unshaded area indicates the feasible region under chance constraints. Figure 3 shows the control norm history for the Monte Carlo simulations. The Monte Carlo simulations disobeyed the chance constraint in $0.01\%$ of total sample size, which meets the prescribed risk bound of $5\%$. Figure 4 plots $x_{T}$ for all samples, overlayed with the terminal covariance constraint ellipsoid $\Sigma_{f}$, sample covariance and the predicted covariance $\Sigma_{T}$ from the optimizer. These numerical results demonstrate the performance of the controller to steer the covariance within the final target value in the given time horizon, while respecting the chance constraints. Moreover, the state trajectories in Figure 2 demonstrate the multi-modal nature of the problem because of the Markovian jump dynamics and mode-dependent controllers. We also note the number of state distribution modes increasing over time, most apparent in Figure 2(a), clearly demonstrating the exponential Gaussian mixture distribution of MJLS, which is hard to control using conventional stochastic optimal control approaches.
@@
Figure 2: State trajectories for Monte Carlo simulations with the state chance constraint $\mathbb{P}(x[2]\geq-10)\geq 0.95$ (shaded region: infeasible.) Both cases use $P(\|u_{k}(i)\|\leq 8)\leq 0.95$.
Figure 3: Control Norm History for the Monte Carlo simulations with state chance constraint $P(x_{2}\geq-10)\geq 0.95$. $u_{pred}(i)$ corresponds to $\|\overline{u}_{k}(i)\|+\sqrt{\tfrac{n_{u}(i)}{\epsilon_{u}}\lambda_{\max}(Y_{k}(i))}$
Figure 4: Sample points from Monte Carlo simulations at final time step, with sampled covariance covariance shown in blue, predicted covariance from optimizer in shaded green, and terminal covariance constraint shown in black. The covariance ellipses are scaled to contain $95\%$ of data of a gaussian distribution with same mean and covariance.
@@ VI Conclusion
In this paper, we consider the optimal covariance steering problem for MJLS under chance constraints. We demonstrate the coupled nature of the problem and decompose the problem into a mean-covariance problem, which can be solved in a two-step convex optimization framework. We propose lossless convex relaxations of the covariance subproblem in both unconstrained and chance-constrained cases and develop an iterative convex optimization framework to solve the chance-constrained covariance steering problem. Numerical simulations for the covariance steering framework are presented, which demonstrate the controller’s performance in steering the covariance of a MJLS to within a given terminal covariance bound while obeying chance constraints with additive noise and bias.
@@ Appendix A MJLS identities
$\mathbb{E}[u_{k}(i)|r_{k}=i]=\frac{\mathbb{E}[u_{k}(i)\mathds{1}_{k}=i]}{\mathbb{E}[\mathds{1}_{r_{k}=i}]}=\frac{\rho_{k}(i)\overline{u}_{k}(i)}{\rho_{k}(i)}=\overline{u}_{k}(i)$ (31)
$\mathbb{E}[(u_{k}(i)-\overline{u}_{k}(i))(u_{k}(i)-\overline{u}_{k}(i))^{\top}|r_{k}=i]$
$=\frac{\mathbb{E}[K_{k}(i)S_{k}(i)K^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]}{\mathbb{E}[\mathds{1}_{r_{k}=i}]}=\frac{Y_{k}(i)}{\rho_{k}(i)}$ (32)
$\mathbb{E}[(x_{k}-\overline{x}_{k}(i))\mathds{1}_{r_{k}=i}]=0$ (33)
$\mathbb{E}[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}]=\mathbb{E}[(x_{k}-\overline{x}_{k}(i)+\overline{x}_{k}(i))(x_{k}-\overline{x}_{k}(i)+$
$\hskip 14.22636pt\overline{x}_{k}(i))^{\top}\mathds{1}_{r_{k}=i}]=S_{k}(i)+0+0+\rho_{k}(i)\overline{x}_{k}(i)\overline{x}^{\top}_{k}(i)$
$=S_{k}(i)+\rho_{k}(i)\overline{x}_{k}(i)\overline{x}^{\top}_{k}(i)$ (34)
$\mathbb{E}[x_{k}u_{k}^{\top}(i)\mathds{1}_{r_{k}=i}]=\mathbb{E}[x_{k}(\overline{u}_{k}(i)+K_{k}(i)$
$\hskip 142.26378pt(x_{k}-\overline{x}_{k}(i)))^{\top}\mathds{1}_{r_{k}=i}]$
$=\mathbb{E}[((x_{k}-\overline{x}_{k}(i))+\overline{x}_{k}(i))(\overline{u}_{k}(i)+$
$\hskip 85.35826ptK_{k}(i)(x_{k}-\overline{x}_{k}(i)))^{\top}\mathds{1}_{r_{k}=i}]$
$=0+S_{k}(i)K^{\top}_{k}(i)+\rho_{k}(i)\overline{x}_{k}(i)\overline{u}^{\top}_{k}(i)+0$
$=\rho_{k}(i)\overline{x}_{k}(i)\overline{u}^{\top}_{k}(i)+S_{k}(i)K^{\top}_{k}(i)$ (35)
$\mathbb{E}[u_{k}(i)c_{k}^{\top}\mathds{1}_{r_{k}=i}]=\mathbb{E}[(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\overline{x}_{k}(i)))c_{k}^{\top}\mathds{1}_{r_{k}=i}]$
@@
$\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]=\mathbb{E}[(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\overline{x}_{k}(i)))(\overline{u}_{k}(i)+$
$\hskip 113.81102ptK_{k}(i)(x_{k}-\overline{x}_{k}(i)))^{\top}\mathds{1}_{r_{k}=i}]$
$=\mathbb{E}[\mathds{1}_{r_{k}=i}]\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+\overline{u}_{k}(i)(\mathbb{E}[x^{\top}_{k}-\overline{x}^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]K_{k}^{\top}(i)$
$\hskip 85.35826pt+K_{k}(i)\mathbb{E}[x_{k}-\overline{x}_{k}(i)\mathds{1}_{r_{k}=i}]\overline{u}_{k}^{\top}(i)$
$\hskip 28.45274pt+K_{k}(i)\mathbb{E}[(x_{k}-\overline{x}_{k}(i))(x_{k}-\overline{x}_{k}(i))^{\top}\mathds{1}_{r_{k}=i}]K_{k}^{\top}(i)$
$=\rho_{k}(i)\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+0+0+K_{k}(i)S_{k}(i)K^{\top}_{k}(i)$
$=\rho_{k}(i)\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+K_{k}(i)S_{k}(i)K_{k}^{\top}(i)$ (37)